# Finite-Sample Analysis of Policy Evaluation for Robust Average Reward Reinforcement Learning

### v1

@@ Abstract
We present the first finite-sample analysis for policy evaluation in robust average-reward Markov Decision Processes (MDPs). Prior works in this setting have established only asymptotic convergence guarantees, leaving open the question of sample complexity. In this work, we address this gap by establishing that the robust Bellman operator is a contraction under the span semi-norm, and developing a stochastic approximation framework with controlled bias. Our approach builds upon Multi-Level Monte Carlo (MLMC) techniques to estimate the robust Bellman operator efficiently. To overcome the infinite expected sample complexity inherent in standard MLMC, we introduce a truncation mechanism based on a geometric distribution, ensuring a finite constant sample complexity while maintaining a small bias that decays exponentially with the truncation level. Our method achieves the order-optimal sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-2})$ for robust policy evaluation and robust average reward estimation, marking a significant advancement in robust reinforcement learning theory.
@@ 1 Introduction
Reinforcement learning (RL) has demonstrated significant success across various domains, including robotics, finance, and healthcare, by enabling agents to learn optimal decision-making strategies through interaction with the environment. However, in many real-world applications, direct interaction with the environment is impractical due to safety concerns, high costs, or data collection constraints (Sünderhauf et al., 2018; Höfer et al., 2021). This challenge is particularly evident in scenarios where agents are trained in simulated environments before being deployed in the real world, such as in robotic control and autonomous driving. The discrepancy between the simulated and real-world environments, known as the simulation-to-reality gap, often leads to performance degradation when the learned policy encounters unmodeled uncertainties. Robust reinforcement learning (robust RL) addresses this challenge by formulating the learning problem as an optimization over an uncertainty set of transition probabilities, ensuring reliable performance under worst-case conditions. In this work, we focus on the problem of evaluating the robust value function and robust average reward for a given policy using only data sampled from a simulator (nominal model), aiming to enhance generalization and mitigate the impact of transition uncertainty in real-world deployment.
@@
Reinforcement learning problems under infinite time horizons are typically studied under two primary reward formulations: the discounted reward setting, where future rewards are exponentially discounted, and the average reward setting, which focuses on optimizing long-term performance. While the discounted-reward formulation is widely used, it may lead to myopic policies that underperform in applications requiring sustained long-term efficiency, such as queueing systems, inventory management, and network control. In contrast, the average-reward setting is more suitable for environments where decisions impact long-term operational efficiency. Despite its advantages, robust reinforcement learning under the average-reward criterion remains largely unexplored. Existing works on robust average-reward RL primarily provide asymptotic guarantees (Wang et al., 2023a, b, 2024), lacking scalable algorithms with finite-time performance bounds. This gap highlights the need for principled approaches that ensure robustness against model uncertainties while maintaining strong long-term performance guarantees.
@@
Solving the robust average-reward reinforcement learning problem is significantly more challenging than its non-robust counterpart, with the primary difficulty arising in policy evaluation. Specifically, the goal is to compute the worst-case value function and worst-case average reward over an entire uncertainty set of transition models while having access only to samples from a nominal transition model. In this paper, we investigate three types of uncertainty sets: Contamination uncertainty sets, TV distance uncertainty sets, and Wasserstein distance uncertainty sets. Unlike the standard average-reward setting, where value functions and average rewards can be estimated directly from observed trajectories, the robust setting introduces an additional layer of complexity due to the need to optimize against adversarial transitions. Consequently, conventional approaches based on direct estimation such as (Wei et al., 2020; Agarwal et al., 2021) immediately fail, as they do not account for the worst-case nature of the problem. Overcoming this challenge requires new algorithmic techniques that can infer the worst-case dynamics using only limited samples from the nominal model.
@@ 1.1 Challenges and Contributions
A common approach to policy evaluation in robust RL is to solve the corresponding robust Bellman operator. However, robust average-reward RL presents additional difficulties compared to the robust discounted setting. In the discounted case, the presence of a discount factor induces a contraction property in the robust Bellman operator (Wang and Zou, 2022; Zhou et al., 2024), facilitating stable iterative updates. In contrast, the average-reward Bellman operator lacks a contraction property with respect to any norm even in the non-robust setting (Zhang et al., 2021), making standard fixed-point analysis inapplicable. Due to this fundamental limitation, existing works on robust average-reward RL such as (Wang et al., 2023b) rely on asymptotic techniques, primarily leveraging ordinary differential equation (ODE) analysis to examine the behavior of temporal difference (TD) learning. These methods exploit the asymptotic stability of the corresponding ODE (Borkar, 2023) to establish almost sure convergence but fail to provide finite-sample performance guarantees. Addressing this limitation requires novel analytical tools and algorithmic techniques capable of providing explicit finite-sample bounds for robust policy evaluation and optimization.
@@
In this work, we first establish and exploit a key structural property of the robust average-reward Bellman operator: its contraction under the span semi-norm, denoted as $\|\cdot\|_{\mathrm{sp}}$. This fundamental result enables the use of stochastic approximation techniques similar to (Zhang et al., 2021) to analyze and bound the error in policy evaluation, overcoming the lack of a standard contraction property that has hindered prior finite-sample analyses. Building on this insight, we develop a novel stochastic approximation framework tailored to the robust average-reward setting. Our approach simultaneously estimates both the robust value function and the robust average reward, leading to an efficient iterative procedure for solving the robust Bellman equation. A critical challenge in this framework uner TV and Wasserstein distance uncertainty sets is accurately estimating the worst-case transition effects, which requires computing the support function of the uncertainty set. While previous work has leveraged Multi-Level Monte Carlo (MLMC) for this task, existing MLMC-based estimators suffer from infinite expected sample complexity due to the unbounded nature of the required geometric sampling. To address this, we introduce a truncation mechanism based on a truncated geometric distribution, ensuring that the sample complexity remains finite while maintaining an exponentially decaying bias. With these techniques, we derive the first finite-sample complexity guarantee for policy evaluation in robust average-reward RL, achieving an optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity bound. The main contributions of this paper are summarized as follows:
@@
  • We prove that under the ergodicity assumption, the robust average-reward Bellman operator is a contraction with respect to the span semi-norm (Theorem 2). This key result enables the application of stochastic approximation techniques for policy evaluation.
@@
  • We prove the convergence of stochastic approximation under the span semi-norm contraction and under i.i.d. with noise with non-zero bias (Theorem 3).
@@
  • We develop an efficient method for computing estimates for the robust Bellman operator across TV distance and Wasserstein distance uncertainty sets. By modifying MLMC with a truncated geometric sampling scheme, we ensure finite expected sample complexity while keeping variance controlled and bias decaying exponentially with truncation level (Theorem 4-6).
@@
2 Related Work
The theoretical gauarantees of robust average reward have been studied by the following works. (Wang et al., 2023a) takes a model-based perspective, approximating robust average-reward MDPs with discounted MDPs and proving uniform convergence of the robust discounted value function as the discount factor approaches one, employing dynamic programming and Blackwell optimality arguments to characterize optimal policies. (Wang et al., 2023b) propose a model-free approach by developing robust relative value iteration (RVI) TD and Q-learning algorithms, proving their almost sure convergence using stochastic approximation, martingale theory, and multi-level Monte Carlo estimators to handle non-linearity in the robust Bellman operator. While these studies provide fundamental insights into robust average-reward RL, they do not establish explicit convergence rate guarantees due to the lack of contraction properties in the robust Bellman operator. In addition, (Sun et al., 2024) studies the policy optimization of average reward robust MDPs with assuming direct queries of the value functions.
Policy evaluation in robust discounted-reward reinforcement learning with finite sample gaurauntees has been extensively studied, with recent works (Wang and Zou, 2022; Zhou et al., 2024; Li et al., 2022; Kumar et al., 2023; Kuang et al., 2022) all focusing on solving the robust Bellman equation by finding its fixed-point solution. This approach is made feasible by the contraction property of the robust Bellman operator under the sup-norm, which arises due to the presence of a discount factor $\gamma<1$. However, this fundamental approach does not directly extend to the robust average-reward setting, where the absence of a discount factor removes the contraction property under any norm. As a result, existing robust discounted methods cannot be applied in the robust average reward RL setting.
@@ 3 Formulation
3.1 Average-reward MDPs.
[…]
A stationary policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps a state to a distribution over $\mathcal{A}$, following which the agent takes action $a$ at state $s$ with probability $\pi(a|s)$. Under a transition kernel $\mathsf{P}$, the average-reward of $\pi$ starting from $s\in\mathcal{S}$ is defined as
The relative value function is defined to measure the cumulative difference between the reward and $g^{\pi}_{\mathsf{P}}$:
$V^{\pi}_{\mathsf{P}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}}\bigg{[}\sum^{\infty}_{t=0}(r_{t}-g^{\pi}_{\mathsf{P}})|S_{0}=s\bigg{]}.$ (2)
Then $(g^{\pi}_{\mathsf{P}},V^{\pi}_{\mathsf{P}})$ satisfies the following Bellman equation (Puterman, 1994):
$V^{\pi}_{\mathsf{P}}(s)=\mathbb{E}_{\pi,\mathsf{P}}\bigg{[}r(s,A)-g^{\pi}_{\mathsf{P}}(s)+\sum_{s^{\prime}\in\mathcal{S}}p^{A}_{s,s^{\prime}}V^{\pi}_{\mathsf{P}}(s^{\prime})\bigg{]}.$ (3)
@@
For robust MDPs, the transition kernel is assumed to be in some uncertainty set $\mathcal{P}$. At each time step, the environment transits to the next state according to an arbitrary transition kernel $\mathsf{P}\in\mathcal{P}$. In this paper, we focus on the $(s,a)$-rectangular compact uncertainty set (Nilim and Ghaoui, 2003; Iyengar, 2005), i.e., $\mathcal{P}=\bigotimes_{s,a}\mathcal{P}^{a}_{s}$, where $\mathcal{P}^{a}_{s}\subseteq\Delta(\mathcal{S})$. Popular uncertainty sets include those defined by the contamination model (Huber, 1965; Wang and Zou, 2022), total variation (Lim et al., 2013), Chi-squared divergence (Iyengar, 2005) and Wasserstein distance (Gao and Kleywegt, 2022).
@@
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{n\geq 0}\mathcal{P}$. It was shown in (Wang et al., 2023a) that the worst case under the time-varying model is equivalent to the one under the stationary model:
@@
Therefore, we limit our focus to the stationary model. We refer to the minimizers of (5) as the worst-case transition kernels for the policy $\pi$, and denote the set of all possible worst-case transition kernels by $\Omega^{\pi}_{g}$, i.e., $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}$.
We focus on the model-free setting, where only samples from the nominal MDP denoted as $\mathsf{P}$ (the centroid of the uncertainty set) are available. We investigate the problem of robust policy evaluation and robust average reward estimation, which means for a given policy $\pi$, we aim to estimate the robust value function and the robust average reward. We now formally define the robust value function $V^{\pi}_{\mathsf{P}_{V}}$ by connecting it with the following robust Bellman equation:
Theorem 1 (Robust Bellman Equation, Theorem 3.1 in (Wang et al., 2023b)).
@@
where $\sigma_{\mathcal{P}^{a}_{s}}(V)=\min_{p\in\mathcal{P}^{a}_{s}}pV$, then the following properties hold:
  1. 1.
    The scalar $g$ corresponds to the robust average reward, i.e., $g=g^{\pi}_{\mathcal{P}}$.
  2. 2.
    The worst-case transition kernel $\mathsf{P}_{V}$ belongs to the set of minimizing transition kernels, i.e., $\mathsf{P}_{V}\in\Omega^{\pi}_{g}$, where
    $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}.$ (7)
  3. 3.
    The function $V$ is unique up to an additive constant, where if $V$ is a solution to the bellman equation, then we have
    $V=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e},$ (8)
    where $c\in\mathbb{R}$ and $\mathbf{e}$ is the all-ones vector in $\mathbb{R}^{|\mathcal{S}|}$.
This robust Bellman equation characterizes the worst-case value function under the uncertainty set. In particular, $\sigma_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$. Unlike the robust discounted case, where the contraction property of the Bellman operator under the sup-norm enables straightforward fixed-point iteration, the robust average-reward Bellman equation does not induce contraction under any norm, making direct iterative methods inapplicable. We now characterize the explicit forms of $\sigma_{\mathcal{P}^{a}_{s}}(V)$ for different compact uncertainty sets are as follows:
Contamination Uncertainty Set The $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\mathsf{P}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Under this uncertainty set, the support function can be computed as
$\sigma_{\mathcal{P}^{a}_{s}}(V)=(1-\delta)\mathsf{P}^{a}_{s}V+\delta\min_{s}V(s),$ (9)
@@
Total Variation Uncertainty Set. The total variation uncertainty set is $\mathcal{P}^{a}_{s}=\{q\in\Delta(|\mathcal{S}|):\frac{1}{2}\|q-\mathsf{P}^{a}_{s}\|_{1}\leq\delta\},$ define $\|\cdot\|_{\mathrm{sp}}$ as the span semi-norm and the support function can be computed using its dual function Iyengar (2005):
$\sigma_{\mathcal{P}^{a}_{s}}(V)$ $=\max_{\mu\geq\mathbf{0}}\big{(}\mathsf{P}^{a}_{s}(V-\mu)-\delta\|V-\mu\|_{\mathrm{sp}}\big{)}.$ (10)
@@
$\mathcal{P}^{a}_{s}=\left\{q\in\Delta(|\mathcal{S}|):W_{l}(\mathsf{P}^{a}_{s},q)\leq\delta\right\}.$ (11)
The support function w.r.t. the Wasserstein distance set, can be calculated as follows (Gao and Kleywegt, 2023):
$\sigma_{\mathcal{P}^{a}_{s}}(V)=\sup_{\lambda\geq 0}\left(-\lambda\delta^{l}+\mathbb{E}_{\mathsf{P}^{a}_{s}}\big{[}\inf_{y}\big{(}V(y)+\lambda d(S,y)^{l}\big{)}\big{]}\right).$ (12)
4 Robust Bellman Operator

### v2

@@ Abstract
We present the first finite-sample analysis of policy evaluation in robust average-reward Markov Decision Processes (MDPs). Prior work in this setting have established only asymptotic convergence guarantees, leaving open the question of sample complexity. In this work, we address this gap by showing that the robust Bellman operator is a contraction under a carefully constructed semi-norm, and developing a stochastic approximation framework with controlled bias. Our approach builds upon Multi-Level Monte Carlo (MLMC) techniques to estimate the robust Bellman operator efficiently. To overcome the infinite expected sample complexity inherent in standard MLMC, we introduce a truncation mechanism based on a geometric distribution, ensuring a finite expected sample complexity while maintaining a small bias that decays exponentially with the truncation level. Our method achieves the order-optimal sample complexity of $\tilde{\mathcal{O}}(\epsilon^{-2})$ for robust policy evaluation and robust average reward estimation, marking a significant advancement in robust reinforcement learning theory.
@@ 1 Introduction
Reinforcement learning (RL) has achieved notable success in domains such as robotics, finance, and healthcare, by enabling agents to learn optimal decision-making strategies through interaction with an environment. However, in many real-world applications, direct interaction is impractical because of safety concerns, high costs, or limited data‑collection budgets Sünderhauf et al. (2018); Höfer et al. (2021). This challenge is particularly evident in scenarios where agents are trained in simulated environments before being deployed in the real world, such as in robotic control and autonomous driving. The mismatch between simulated and real environments, known as the simulation-to-reality gap, often leads to performance degradation when the learned policy encounters unmodeled uncertainties. Robust reinforcement learning (robust RL) addresses this challenge by formulating the learning problem as an optimization over an uncertainty set of transition probabilities, ensuring reliable performance under worst-case conditions. In this work, we focus on the problem of evaluating the robust value function and robust average reward for a given policy using only data sampled from a simulator (nominal model), aiming to enhance generalization and mitigate the impact of transition uncertainty in real-world deployment.
@@
Reinforcement learning problems under infinite time horizons are typically studied under two primary reward formulations: the discounted reward setting, where future rewards are exponentially discounted, and the average reward setting, which focuses on optimizing long-term performance. While the discounted-reward formulation is widely used, it may lead to myopic policies that underperform in applications requiring sustained long-term efficiency, such as queueing systems, inventory management, and network control. In contrast, the average-reward setting is more suitable for environments where decisions impact long-term operational efficiency. Despite its advantages, robust reinforcement learning under the average-reward criterion remains largely unexplored. Existing works on robust average-reward RL primarily provide asymptotic guarantees Wang et al. (2023a, b, 2024), lacking algorithms with finite-time performance bounds. This gap highlights the need for principled approaches that ensure robustness against model uncertainties while maintaining strong long-term performance guarantees.
@@
Solving the robust average-reward reinforcement learning problem is significantly more challenging than its non-robust counterpart, with the primary difficulty arising in policy evaluation. Specifically, the goal is to compute the worst-case value function and worst-case average reward over an entire uncertainty set of transition models while having access only to samples from a nominal transition model. In this paper, we investigate three types of uncertainty sets: Contamination uncertainty sets, TV distance uncertainty sets, and Wasserstein distance uncertainty sets. Unlike the standard average-reward setting, where value functions and average rewards can be estimated directly from observed trajectories, the robust setting introduces an additional layer of complexity due to the need to optimize against adversarial transitions. Consequently, conventional approaches based on direct estimation such as Wei et al. (2020); Agarwal et al. (2021) immediately fail, as they do not account for the worst-case nature of the problem. Overcoming this challenge requires new algorithmic techniques that can infer the worst-case dynamics using only limited samples from the nominal model.
@@ 1.1 Challenges and Contributions
A common approach to policy evaluation in robust RL is to solve the corresponding robust Bellman operator. However, robust average-reward RL presents additional difficulties compared to the robust discounted setting. In the discounted case, the presence of a discount factor induces a contraction property in the robust Bellman operator Wang and Zou (2022); Zhou et al. (2024), facilitating stable iterative updates. In contrast, the average-reward Bellman operator lacks a contraction property with respect to any norm even in the non-robust setting Zhang et al. (2021), making standard fixed-point analysis inapplicable. Due to this fundamental limitation, existing works on robust average-reward RL such as Wang et al. (2023b) rely on asymptotic techniques, primarily leveraging ordinary differential equation (ODE) analysis to examine the behavior of temporal difference (TD) learning. These methods exploit the asymptotic stability of the corresponding ODE Borkar (2023) to establish almost sure convergence but fail to provide finite-sample performance guarantees. Addressing this limitation requires novel analytical tools and algorithmic techniques capable of providing explicit finite-sample bounds for robust policy evaluation and optimization.
In this work, we first establish and exploit a key structural property of the robust average-reward Bellman operator with uncertainty set $\mathcal{P}$ under the ergodicity assumption: its contraction under some semi-norm, denoted as $\|\cdot\|_{\mathcal{P}}$, where the detailed construction is specified in Theorem 2 and (15). Constructing $\|\cdot\|_{\mathcal{P}}$ is not straightforward, because ergodicity alone only guarantees that the chain mixes over multiple steps and fails to produce a single‐step contraction for familiar measures such as the span semi-norm. To overcome this, we group together all the worst‐case transition dynamics under uncertainty into one compact family of linear mappings, and observe that their “worst‐case gain” over any number of steps stays strictly below $1$. From this we build an extremal norm, which by construction shrinks every non‐constant component by the same fixed factor in a single step. Finally, we add a small “quotient” correction that exactly annihilates constant shifts, producing a semi-norm that vanishes only on constant functions but still inherits the one‐step shrinkage. The above construction yields a uniform, strict contraction for the robust Bellman operator.
@@
This fundamental result above enables the use of stochastic approximation techniques similar to Zhang et al. (2021) to analyze and bound the error in policy evaluation, overcoming the lack of a standard contraction property that has hindered prior finite-sample analyses. Building on this insight, we develop a novel stochastic approximation framework tailored to the robust average-reward setting. Our approach simultaneously estimates both the robust value function and the robust average reward, leading to an efficient iterative procedure for solving the robust Bellman equation. A critical challenge in this framework under TV and Wasserstein distance uncertainty sets is accurately estimating the worst-case transition effects, which requires computing the support function of the uncertainty set. While previous works Blanchet and Glynn (2015); Blanchet et al. (2019); Wang et al. (2023b) have leveraged Multi-Level Monte Carlo (MLMC) for this task, their MLMC-based estimators suffer from infinite expected sample complexity due to the unbounded nature of the required geometric sampling, which lead to only asymptotic convergence. To address this, we introduce a truncation mechanism based on a truncated geometric distribution, ensuring that the sample complexity remains finite while maintaining an exponentially decaying bias. With these techniques, we derive the first finite-sample complexity guarantee for policy evaluation in robust average-reward RL, achieving an optimal $\tilde{\mathcal{O}}(\epsilon^{-2})$ sample complexity bound. The main contributions of this paper are summarized as follows:
@@
  • We prove that under the ergodicity assumption, the robust average-reward Bellman operator is a contraction with respect to a suitably constructed semi-norm (Theorem 2). This key result enables the application of stochastic approximation techniques for policy evaluation.
@@
  • We prove the convergence of stochastic approximation under the semi-norm contraction and under i.i.d. with noise with non-zero bias (Theorem 8) as an intermediate result.
@@
  • We develop an efficient method for computing estimates for the robust Bellman operator under TV distance and Wasserstein distance uncertainty sets. By modifying MLMC with a truncated geometric sampling scheme, we ensure finite expected sample complexity while keeping variance controlled and bias decaying exponentially with truncation level (Theorem 3-5).
@@ 2 Formulation
For a robust MDP with state space $\mathcal{S}$ and action space $\mathcal{A}$ while $|\mathcal{S}|=S$ and $|\mathcal{A}|=A$, the transition kernel is assumed to be in some uncertainty set $\mathcal{P}$. At each time step, the environment transits to the next state according to an arbitrary transition kernel $\mathsf{P}\in\mathcal{P}$. In this paper, we focus on the $(s,a)$-rectangular compact uncertainty set Nilim and Ghaoui (2003); Iyengar (2005), i.e., $\mathcal{P}=\bigotimes_{s,a}\mathcal{P}^{a}_{s}$, where $\mathcal{P}^{a}_{s}\subseteq\Delta(\mathcal{S})$. Popular uncertainty sets include those defined by the contamination model Huber (1965); Wang and Zou (2022), total variation Lim et al. (2013), and Wasserstein distance Gao and Kleywegt (2022).
@@
where $\kappa=(\mathsf{P}_{0},\mathsf{P}_{1}...)\in\bigotimes_{n\geq 0}\mathcal{P}$. It was shown in Wang et al. (2023a) that the worst case under the time-varying model is equivalent to the one under the stationary model:
@@
Therefore, we limit our focus to the stationary model. We refer to the minimizers of (2) as the worst-case transition kernels for the policy $\pi$, and denote the set of all possible worst-case transition kernels by $\Omega^{\pi}_{g}$, i.e., $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}$, where $g^{\pi}_{\mathsf{P}}$ denotes the average reward of policy $\pi$ under the single transition $\mathsf{P}\in\mathcal{P}$:
We focus on the model-free setting, where only samples from the nominal MDP denoted as $\tilde{\mathsf{P}}$ (the centroid of the uncertainty set) are available. We investigate the problem of robust policy evaluation and robust average reward estimation, which means for a given policy $\pi$, we aim to estimate the robust value function and the robust average reward. We now formally define the robust value function $V^{\pi}_{\mathsf{P}_{V}}$ by connecting it with the following robust Bellman equation:
Theorem 1 (Robust Bellman Equation, Theorem 3.1 in Wang et al. (2023b)).
@@
where $\sigma_{\mathcal{P}^{a}_{s}}(V)=\min_{p\in\mathcal{P}^{a}_{s}}pV$ is denoted as the support function, then the scalar $g$ corresponds to the robust average reward, i.e., $g=g^{\pi}_{\mathcal{P}}$, and the worst-case transition kernel $\mathsf{P}_{V}$ belongs to the set of minimizing transition kernels, i.e., $\mathsf{P}_{V}\in\Omega^{\pi}_{g}$, where $\Omega^{\pi}_{g}\triangleq\{\mathsf{P}\in\mathcal{P}:g^{\pi}_{\mathsf{P}}=g^{\pi}_{\mathcal{P}}\}$. Furthermore, the function $V$ is unique up to an additive constant, where if $V$ is a solution to the Bellman equation, then we have $V=V^{\pi}_{\mathsf{P}_{V}}+c\mathbf{e}$, where $c\in\mathbb{R}$ and $\mathbf{e}$ is the all-ones vector in $\mathbb{R}^{S}$, and $V^{\pi}_{\mathsf{P}_{V}}$ is definied as the relative value function of the policy $\pi$ under the single transition $\mathsf{P}_{V}$ as follows:
$V^{\pi}_{\mathsf{P}_{V}}(s)\triangleq\mathbb{E}_{\pi,\mathsf{P}_{V}}\bigg{[}\sum^{\infty}_{t=0}(r_{t}-g^{\pi}_{\mathsf{P}_{V}})|S_{0}=s\bigg{]}.$ (5)
Theorem 1 implies that the robust Bellman equation (4) identifies both the worst‐case average reward $g$ and a corresponding value function $V$ that is determined only up to an additive constant. In particular, $\sigma_{\mathcal{P}^{a}_{s}}(V)$ represents the worst-case transition effect over the uncertainty set $\mathcal{P}^{a}_{s}$. Unlike the robust discounted case, where the contraction property of the Bellman operator under the sup-norm enables straightforward fixed-point iteration, the robust average-reward Bellman equation does not induce contraction under any norm, making direct iterative methods inapplicable. Throughout the paper, we denote $\mathbf{e}$ as the all-ones vector in $\mathbb{R}^{S}$. We now characterize the explicit forms of $\sigma_{\mathcal{P}^{a}_{s}}(V)$ for different compact uncertainty sets are as follows:
Contamination Uncertainty Set The $\delta$-contamination uncertainty set is $\mathcal{P}^{a}_{s}=\{(1-\delta)\tilde{\mathsf{P}}^{a}_{s}+\delta q:q\in\Delta(\mathcal{S})\},$ where $10<\delta<$ 1 is the radius. Under this uncertainty set, the support function can be computed as
$\sigma_{\mathcal{P}^{a}_{s}}(V)=(1-\delta)\tilde{\mathsf{P}}^{a}_{s}V+\delta\min_{s}V(s),$ (6)
@@
Total Variation Uncertainty Set. The total variation uncertainty set is $\mathcal{P}^{a}_{s}=\{q\in\Delta(|\mathcal{S}|):\frac{1}{2}\|q-\tilde{\mathsf{P}}^{a}_{s}\|_{1}\leq\delta\},$ define $\|\cdot\|_{\mathrm{sp}}$ as the span semi-norm and the support function can be computed using its dual function Iyengar (2005):
$\sigma_{\mathcal{P}^{a}_{s}}(V)=\max_{\mu\geq\mathbf{0}}\big{(}\tilde{\mathsf{P}}^{a}_{s}(V-\mu)-\delta\|V-\mu\|_{\mathrm{sp}}\big{)}.$ (7)
@@
$\mathcal{P}^{a}_{s}=\left\{q\in\Delta(|\mathcal{S}|):W_{l}(\tilde{\mathsf{P}}^{a}_{s},q)\leq\delta\right\}.$ (8)
The support function w.r.t. the Wasserstein distance set, can be calculated as follows Gao and Kleywegt (2023):
$\sigma_{\mathcal{P}^{a}_{s}}(V)=\sup_{\lambda\geq 0}\left(-\lambda\delta^{l}+\mathbb{E}_{\tilde{\mathsf{P}}^{a}_{s}}\big{[}\inf_{y}\big{(}V(y)+\lambda d(S,y)^{l}\big{)}\big{]}\right).$ (9)
2.2 Robust Bellman Operator